THe height in feet a ball is t seconds after being throan in the air is given by H(t)=-16t^2+64t+5
A)How long did it take the ball to reach its maximum height above the ground
B)What was the maximum height above the ground the ball reached

Answer by nerdybill(7384)

THe height in feet a ball is t seconds after being throan in the air is given by H(t)=-16t^2+64t+5
A)How long did it take the ball to reach its maximum height above the ground
Vertex of parabola is the max:
t = -b/(2a)
t = -64/(2(-16))
t = -64/(-32)
t = 2 seconds
.
B)What was the maximum height above the ground the ball reached
plug t=2 into:
H(t)=-16t^2+64t+5
H(2)=-16(2)^2+64(2)+5
H(2)=-16(4)+128+5
H(2)=-64+128+5
H(2)= 69 feet